# Tuning spin-density separation via finite-range interactions: Dimensionality-driven signatures in dynamic structure factors

### v1

@@ Abstract
Spin-density separation, marked by distinct propagation velocities of spin and density excitations, epitomizes strong correlations, historically confined to one-dimensional (1D) systems. Recent ultracold quantum gas experiments, however, demonstrate its emergence in higher dimensions through precise tuning of intra- to interspecies interaction ratios, inspiring exploration of how dimensionality and interatomic interactions govern quantum correlations. We investigate this in two-component bosonic mixtures with finite-range interactions, probing 1D and 3D dynamics. Using effective field theory within the one-loop approximation, we derive analytical expressions for zero-temperature ground-state energy and quantum depletion, seamlessly recovering contact interaction results in the contact limit. By crafting an effective action for decoupled density and spin modes, we compute dynamic structure factors (DSFs), revealing how finite-range interactions sculpt spin-density separation. A pivotal finding is the dimensionality-driven divergence in DSF peak dynamics: in 1D, peaks ascend to higher frequencies with increasing interaction strength, yielding sharp responses; in 3D, peaks descend to lower frequencies, with broader density wave profiles. These insights highlight dimensionality’s critical role in collective excitations and provide a robust theoretical blueprint for probing interaction-driven quantum phenomena via Bragg spectroscopy, paving new pathways for exploring dimensionally tuned quantum correlations in ultracold quantum gases.
@@ I INTRODUCTION
Spin-density separation, a defining feature of one-dimensional (1D) quantum many-body systems, emerges as the decoupling of low-energy collective excitations into independent spin and density modes, elegantly captured by the Tomonaga-Luttinger liquid framework for 1D Fermi gases [1, 2, 3, 4, 5]. Governed by tunable interactions, this phenomenon results in distinct linear dispersions for spin and density excitations, with velocities determined by Luttinger parameters and interaction strength [6, 7, 8, 9, 10, 11]. Experimental techniques, such as Bragg spectroscopy [12, 13] and tunneling spectroscopy [14, 15], probe the dynamic structure factor (DSF) and tunneling conductance of one-dimensional quantum systems, revealing distinct spectral features of spin and density excitations, which align with theoretical predictions and highlight the role of interaction strength in shaping excitation dynamics [16]. Although primarily observed in 1D Fermi systems, spin-density separation extends to 1D Bosonic mixtures, suggesting its potential in higher-dimensional bosonic systems [17]. However, whether spin-density separation persists in higher-dimensional systems remains an important question, reflecting the critical role of dimensionality in modulating quantum correlations and collective phenomena.
@@
Another important motivation for this work stems from the emergence of exotic many-body phenomena—including quantum droplets [20, 21, 22], phase separation [23, 24, 25], mixed bubble states [26, 27], and miscible-immiscible transitions [28, 29, 30]—in two-component Bosonic mixtures. These phenomena are driven by the intricate interplay between interactions and quantum fluctuations. In the weakly interacting regime, they are predominantly governed by the competition between intra- and interspecies $s$-wave contact interactions, where the $s$-wave scattering length serves as the primary tunable parameter [31, 32, 33]. Beyond this standard paradigm, a diverse array of interactions—such as dipole-dipole interactions, Rabi coupling, spin-orbit coupling, and finite-range interactions—introduce additional degrees of freedom. These include non-locality, spin coherence, and momentum-dependent coupling, enabling precise sculpting of the effective interatomic potential [34, 35, 36, 37, 38, 39, 40]. Such interactions significantly expand the spectrum of accessible quantum phases. Crucially, finite-range interactions have recently emerged as a critically important control parameter [41, 42, 43, 44, 45, 46]. In two-component systems, they play a pivotal role in stabilizing quantum droplets and modulating density profiles beyond the contact-interaction approximation [47]. Furthermore, incorporating intra- and interspecies effective ranges ($r_{\text{e}}$, $r_{\text{e}_{12}}$) facilitates independent control over both density and spin channels, enabling comprehensive interaction-driven modulation of collective excitations.
@@ II.1 Partition function of the system
[…]
@@
In Eq. (2), the $\Psi\left(\bf{r},\tau\right)=[\psi_{1},\psi_{2}]^{T}$ represent the two-component atomic bosonic complex fields, varying in both space $\bf{r}$ and imaginary time $\tau$. The parameters $\mu$ and $\beta=(k_{\text{B}}T)^{-1}$ respectively define the thermodynamic chemical potential and the inverse thermal energy scale governing system equilibrium. The $g^{(0)}$ and $g^{(0)}_{12}$ denote the intra- and interspecies zero-range interatomic interaction coupling constants, respectively, with $g^{(0)}\neq g^{(0)}_{12}$ in view of the relevant experiment. Crucially, the higher-order terms $g^{(2)}$ and $g^{(2)}_{12}$ in Eq. (2) encode finite-range interaction effects through the $s$-wave pseudopotential expansion, where $\mathcal{O}(k^{2})$ momentum dependence induces beyond-mean-field correlations. The explicit forms of the interaction parameters $g^{(0)}$ and $g^{(2)}$ will be systematically derived in the subsequent analysis, establishing their competition as the microscopic origin of spin-density separation in the phase diagram.
@@
where $a^{\left(\text{1D}\right)}$ and $a_{12}^{\left(\text{1D}\right)}$ are the effective 1D scattering lengths, and $C=-\zeta(1/2)/\sqrt{2}\simeq 1.0326$ with $\zeta(x)$ being the Riemann zeta function. For strong confinement ($l_{\perp}\gg a^{\left(\text{1D}\right)},a_{12}^{\left(\text{1D}\right)}$), these simplify to $g^{\left(0\right)\left(\text{1D}\right)}\approx 2\hbar^{2}a^{\left(\text{1D}\right)}/ml_{\perp}^{2}$ and $g^{\left(0\right)\left(\text{1D}\right)}_{12}\approx 2\hbar^{2}a_{12}^{\left(\text{1D}\right)}/ml_{\perp}^{2}$ [50]. The finite-range interaction strengths are assumed to be $g^{\left(2\right)\left(\text{1D}\right)}=-\hbar^{2}r_{\text{e}}^{\left(\text{1D}\right)}/m$ and $g^{\left(2\right)\left(\text{1D}\right)}_{12}=-\hbar^{2}r_{\text{e}_{12}}^{\left(\text{1D}\right)}/m$, where $r_{\text{e}}^{\left(\text{1D}\right)}$ and $r_{\text{e}_{12}}^{\left(\text{1D}\right)}$ are the intra- and interspecies $s$-wave effective 1D ranges.
@@
Now, we are ready to investigate the non-trivial interplay between dimensional confinement and the microscopic interaction structure by calculating both the equation of state (EOS) and the DSF. In this context, we limit ourselves to the case that intraspecies interactions are all repulsive ($g^{\left(0\right)}>0$, $g^{\left(2\right)}>0$) and we work in the superfluid phase, where a U(1) gauge symmetry of each component is spontaneously broken [35]. Under this assumption, the complex fields in Eq. (3) can be expanded as $\psi_{i}\left({\bf r},\tau\right)=\sqrt{n_{0}}+\eta_{i}\left({\bf r},\tau\right)$, where $n_{0}=|\psi^{\left(0\right)}_{i}|^{2}$ represents the 3D condensate or 1D quasi-condensate density within the Bogoliubov approximation [48, 51], while $\eta_{i}^{*}\left({\bf r},\tau\right)$ and $\eta_{i}\left({\bf r},\tau\right)$ denote the fluctuation fields above the condensate. The mean-field plus Gaussian approximation is obtained by expanding Eq. (3) up to second order in $\eta_{i}^{*}\left({\bf r},\tau\right)$ and $\eta_{i}\left({\bf r},\tau\right)$.
@@
Figure 1: (a) Dimensionless excitation spectrum $f_{+}(k^{\prime})$ from Eq. (19) for varying $z_{\rho}\equiv 4mng_{\rho}^{(2)}/\hbar^{2}$. (b) Corresponding spectrum $f_{-}(k^{\prime})$ for different $z_{\sigma}\equiv 4mng_{\sigma}^{(2)}/\hbar^{2}$. The dimensionless interaction parameter is fixed at $y=0.3$.
@@ II.2 Bogoliubov excitations
$-\hbar G_{B\bm{k}n}^{-1}$ $=$ $\left(\begin{array}[]{cc}-i\hbar\omega_{n}+\epsilon^{0}_{k}+g^{\prime}n_{0},&g^{\prime}n_{0},\\ g^{\prime}n_{0},&i\hbar\omega_{n}+\epsilon^{0}_{k}+g^{\prime}n_{0},\end{array}\right),$ (11)
with $\epsilon^{0}_{k}=\hbar^{2}k^{2}/2m$ and $g^{\prime}=g^{\left(0\right)}+g^{\left(2\right)}k^{2}$. The off-diagonal matrix in Eq. (8), representing the self-energy due to interspecies coupling, is expressed as $\hbar\Sigma_{12}=\left(g^{\left(0\right)}_{12}+g^{\left(2\right)}_{12}k^{2}\right)n_{0}I$.
@@
$\bm{M}\!\!$ =\!\!= $\left(\begin{array}[]{cccc}\!\!\epsilon^{0}_{k}+g^{\prime}n_{0}\!&g^{\prime}n_{0}&g^{\prime}_{12}n_{0}&g^{\prime}_{12}n_{0}\\ g^{\prime}n_{0}&\!\!\epsilon^{0}_{k}+g^{\prime}n_{0}\!&g^{\prime}_{12}n_{0}&g^{\prime}_{12}n_{0}\\ g^{\prime}_{12}n_{0}&g^{\prime}_{12}n_{0}&\!\!\epsilon^{0}_{k}+g^{\prime}n_{0}\!&g^{\prime}n_{0}\\ g^{\prime}_{12}n_{0}&g^{\prime}_{12}n_{0}&g^{\prime}n_{0}&\!\!\epsilon^{0}_{k}+g^{\prime}n_{0}\!\end{array}\right),$ (17)
@@
Figure 1 shows excitation spectra $f_{\pm}$ versus dimensionless wave vector $k^{\prime}=\hbar k/\sqrt{mg^{(0)}n_{0}}$ for various finite-range couplings $z_{\rho(\sigma)}$. We have checked that Eq. (19) is simplified into Ref. [53] for vanishing the finite-range interaction coupling constant $z_{\rho(\sigma)}=0$. The competition between contact ($y$) and finite-range ($z$) interactions distinctly modifies density and spin excitations in Figs. 1(a) and (b), respectively. In the phonon regime ($k\to 0$):
@@
Spinor $y<-1$, $z_{\rho(\sigma)}=0$ Quantum droplet [22]
$y<-1$, $z_{\rho(\sigma)}>0$ Finite-range quantum droplet [37]
$10<y<$ 1, $z_{\rho(\sigma)}=0$ Spin-density separation [18]
$10<y<$ 1, $0\leq z_{\rho(\sigma)}\leq 1$ Spin-density separation (finite-range) This work
@@
(i) Single-component system ($y=0$): Finite-range interactions ($z_{\rho(\sigma)}\neq 0$) govern the EOS [41, 42, 45, 43]. Positive $z_{\rho(\sigma)}>0$ induces a dimensional crossover [44, 46], while $z_{\rho(\sigma)}=0$ yields a distinct dimensional crossover [54, 55, 56].
@@
$z_{\rho(\sigma)}=0$: Quantum droplet phase for $y<-1$ (stabilized by beyond-mean-field effects [22]) and $-1<y<0$ [22]; spin-density separation for $10<y<$ 1 [18].
$z_{\rho(\sigma)}\neq 0$: Quantum droplets persist for $y<-1$ [37]. For $10<y<$ 1 and $0\leq z_{\rho(\sigma)}\leq 1$, finite-range interactions modify spin-density separation (this work’s focus). The system transitions to a mixed bubble phase at $y>0$ [26] and phase separation at $y>1$ [57].
@@
Figure 2: (a) Scaling function $h(z_{\rho})$ in Eq. (25) for $z_{\rho}\equiv 4mng_{\rho}^{(2)}/\hbar^{2}$ varied in $[0,0.7]$ and discrete values of $z_{\sigma}\equiv 4mng_{\sigma}^{(2)}/\hbar^{2}$. (b) Scaling function $I(z_{\sigma})$ in Eq. (29) for $z_{\sigma}\equiv 4mng_{\sigma}^{(2)}/\hbar^{2}$ varied in $[0,0.7]$ and discrete values of $z_{\rho}$.
III EQUATION OF STATE IN $d$-DIMENSIONAL SYSTEM

### v2

@@ Abstract
Spin-density (charge) separation, marked by distinct propagation velocities of spin and density excitations, epitomizes strong correlations, historically confined to one-dimensional (1D) systems. The recent experimental work of S. Dhar, B. Wang, M. Horvath, et al. Nature 642, 53 (2025), using a weakly interacting 3D Bose-Einstein condensate of 133Cs atoms confined in a 2D optical lattice to realize spin-density separation and demonstrate boson anyonization, motivates a deeper exploration into how dimensionality and interactions govern quantum correlations. In this work, we investigate this in two-component bosonic mixtures with finite-range interactions, probing 1D and 3D dynamics. Using path integral effective field theory within the one-loop approximation, we derive analytical expressions for zero-temperature ground-state energy and quantum depletion, seamlessly recovering contact interaction results in the contact limit. By crafting an effective action for decoupled density and spin modes, we compute dynamic structure factors (DSFs), revealing how finite-range interactions sculpt spin-density separation. A pivotal finding is the dimensionality-driven divergence in DSF peak dynamics: in 1D, peaks ascend to higher frequencies with increasing interaction strength, yielding sharp responses; in 3D, peaks descend to lower frequencies, with broader density wave profiles. These insights highlight dimensionality’s critical role in collective excitations and provide a robust theoretical blueprint for probing interaction-driven quantum phenomena via Bragg spectroscopy, paving new pathways for exploring dimensionally tuned quantum correlations in ultracold quantum gases.
Ultracold quantum gas, Effective field theory, Dynamical structure factor, Quantum depletion, Spin-density (charge) separation
@@ I INTRODUCTION
Spin-density separation, a defining feature of one-dimensional (1D) quantum many-body systems, emerges as the decoupling of low-energy collective excitations into independent spin and density modes, elegantly captured by the Tomonaga-Luttinger liquid framework for 1D Fermi gases [1, 2, 3, 4, 5]. Governed by tunable interactions, this phenomenon results in distinct linear dispersions for spin and density excitations, with velocities determined by Luttinger parameters and interaction strength [6, 7, 8, 9, 10, 11]. Experimental techniques, such as Bragg spectroscopy [12, 13] and tunneling spectroscopy [14, 15], probe the dynamic structure factor (DSF) and tunneling conductance of one-dimensional quantum systems, revealing distinct spectral features of spin and density excitations, which align with theoretical predictions and highlight the role of interaction strength in shaping excitation dynamics [16]. Although primarily observed in 1D Fermi systems, spin-density separation extends to 1D Bosonic mixtures, suggesting its potential in higher-dimensional bosonic systems [17]. In a 1D bosonic gas, spin-charge separation enables the experimental observation of anyonic correlations, highlighting tunable statistical phases via strong interactions [18]. However, whether spin-density separation persists in higher-dimensional systems remains an important question, reflecting the critical role of dimensionality in modulating quantum correlations and collective phenomena.
@@
Another important motivation for this work stems from the emergence of exotic many-body phenomena—including quantum droplets [21, 22, 23], phase separation [24, 25, 26], mixed bubble states [27, 28], and miscible-immiscible transitions [29, 30, 31]—in two-component Bosonic mixtures. These phenomena are driven by the intricate interplay between interactions and quantum fluctuations. In the weakly interacting regime, they are predominantly governed by the competition between intra- and interspecies $s$-wave contact interactions, where the $s$-wave scattering length serves as the primary tunable parameter [32, 33, 34]. Beyond this standard paradigm, a diverse array of interactions—such as dipole-dipole interactions, Rabi coupling, spin-orbit coupling, and finite-range interactions—introduce additional degrees of freedom. These include non-locality, spin coherence, and momentum-dependent coupling, enabling precise sculpting of the effective interatomic potential [35, 36, 37, 38, 39, 40, 41, 42]. Such interactions significantly expand the spectrum of accessible quantum phases. Crucially, finite-range interactions have recently emerged as a critically important control parameter [43, 44, 45, 46, 47, 48]. In two-component systems, they play a pivotal role in stabilizing quantum droplets and modulating density profiles beyond the contact-interaction approximation [49, 50]. Furthermore, incorporating intra- and interspecies effective ranges ($r_{\text{e}}$, $r_{\text{e}_{12}}$) facilitates independent control over both density and spin channels, enabling comprehensive interaction-driven modulation of collective excitations.
@@ II.1 Partition function of the system
with the action functional in Eq. (1) reading $S\left[\bar{\Psi},\Psi\right]=\int_{0}^{\hbar\beta}\mathrm{d}\tau\int_{L^{d}}\mathrm{d}^{d}\mathbf{r}\penalty 10000\ \mathcal{L}\left[\bar{\Psi},\Psi\right]$. The concrete expression of Lagrangian density $\mathcal{L}\left[\bar{\Psi},\Psi\right]$ here takes the form
@@
In Eq. (2), the $\Psi\left(\bm{r},\tau\right)=[\psi_{1},\psi_{2}]^{\mathrm{T}}$ represent the two-component atomic bosonic complex fields, varying in both space $\bm{r}$ and imaginary time $\tau$. The parameters $\mu$ and $\beta=(k_{\text{B}}T)^{-1}$ respectively define the thermodynamic chemical potential and the inverse thermal energy scale governing system equilibrium. The $g^{(0)}$ and $g^{(0)}_{12}$ denote the intra- and interspecies zero-range interatomic interaction coupling constants, respectively, with $g^{(0)}\neq g^{(0)}_{12}$ in view of the relevant experiment. Crucially, the higher-order terms $g^{(2)}$ and $g^{(2)}_{12}$ in Eq. (2) encode finite-range interaction effects through the $s$-wave pseudopotential expansion, where $\mathcal{O}(k^{2})$ momentum dependence induces beyond-mean-field correlations. The explicit forms of the interaction parameters $g^{(0)}$ and $g^{(2)}$ will be systematically derived in the subsequent analysis, establishing their competition as the microscopic origin of spin-density separation in the phase diagram.
@@
where $a^{\left(\text{1D}\right)}$ and $a_{12}^{\left(\text{1D}\right)}$ are the effective 1D scattering lengths, and $C=-\zeta(1/2)/\sqrt{2}\simeq 1.0326$ with $\zeta(x)$ being the Riemann zeta function [53]. For strong confinement ($l_{\perp}\gg a^{\left(\text{1D}\right)},a_{12}^{\left(\text{1D}\right)}$), these simplify to $g^{\left(0\right)\left(\text{1D}\right)}\approx 2\hbar^{2}a^{\left(\text{1D}\right)}/ml_{\perp}^{2}$ and $g^{\left(0\right)\left(\text{1D}\right)}_{12}\approx 2\hbar^{2}a_{12}^{\left(\text{1D}\right)}/ml_{\perp}^{2}$ [54]. The finite-range interaction strengths are assumed to be $g^{\left(2\right)\left(\text{1D}\right)}=-\hbar^{2}r_{\text{e}}^{\left(\text{1D}\right)}/m$ and $g^{\left(2\right)\left(\text{1D}\right)}_{12}=-\hbar^{2}r_{\text{e}_{12}}^{\left(\text{1D}\right)}/m$, where $r_{\text{e}}^{\left(\text{1D}\right)}$ and $r_{\text{e}_{12}}^{\left(\text{1D}\right)}$ are the intra- and interspecies $s$-wave effective 1D ranges.
@@
Now, we are ready to investigate the non-trivial interplay between dimensional confinement and the microscopic interaction structure by calculating both the ground-state energy, quantum depletion and the DSF. In this context, we limit ourselves to the case that intraspecies interactions are all repulsive ($g^{\left(0\right)}>0$, $g^{\left(2\right)}>0$) and we work in the superfluid phase, where a U(1) gauge symmetry of each component is spontaneously broken [36]. Under this assumption, the complex fields in Eq. (3) can be expanded as $\psi_{i}\left({\bm{r}},\tau\right)=\sqrt{n_{0}}+\eta_{i}\left({\bm{r}},\tau\right)$, where $n_{0}=|\psi^{\left(0\right)}_{i}|^{2}$ represents the 3D condensate or 1D quasi-condensate density within the Bogoliubov approximation [51, 53], while $\eta_{i}^{*}\left({\bm{r}},\tau\right)$ and $\eta_{i}\left({\bm{r}},\tau\right)$ denote the fluctuation fields above the condensate. The mean-field plus Gaussian approximation is obtained by expanding Eq. (3) up to second order in $\eta_{i}^{*}\left({\bm{r}},\tau\right)$ and $\eta_{i}\left({\bm{r}},\tau\right)$.
@@
Figure 1: (a) Dimensionless excitation spectrum $f_{+}(k^{\prime})$ from Eq. (19) for varying $z_{\rho}\equiv 4mn_{0}g_{\rho}^{(2)}/\hbar^{2}$. (b) Corresponding spectrum $f_{-}(k^{\prime})$ for different $z_{\sigma}\equiv 4mn_{0}g_{\sigma}^{(2)}/\hbar^{2}$. The dimensionless interaction parameter is fixed at $y=0.3$.
@@ II.2 Bogoliubov excitations
$-\hbar G_{B\bm{k}n}^{-1}$ $=$ $\left(\begin{array}[]{cc}-\mathrm{i}\hbar\omega_{n}+\epsilon^{0}_{k}+g^{\prime}n_{0},&g^{\prime}n_{0},\\ g^{\prime}n_{0},&\mathrm{i}\hbar\omega_{n}+\epsilon^{0}_{k}+g^{\prime}n_{0},\end{array}\right),$ (11)
with $\epsilon^{0}_{k}=\hbar^{2}k^{2}/2m$ and $g^{\prime}=g^{\left(0\right)}+g^{\left(2\right)}k^{2}$. The off-diagonal matrix in Eq. (8), representing the self-energy due to interspecies coupling, is expressed as $\hbar\Sigma_{12}=\left(g^{\left(0\right)}_{12}+g^{\left(2\right)}_{12}k^{2}\right)n_{0}I$, where $I$ is a $2\times 2$ matrix with all elements equal to $1$.
@@
$\bm{M}\!\!$ $=\!\!$ $\left(\begin{array}[]{cccc}\!\!\epsilon^{0}_{k}+g^{\prime}n_{0}\!&g^{\prime}n_{0}&g^{\prime}_{12}n_{0}&g^{\prime}_{12}n_{0}\\ g^{\prime}n_{0}&\!\!\epsilon^{0}_{k}+g^{\prime}n_{0}\!&g^{\prime}_{12}n_{0}&g^{\prime}_{12}n_{0}\\ g^{\prime}_{12}n_{0}&g^{\prime}_{12}n_{0}&\!\!\epsilon^{0}_{k}+g^{\prime}n_{0}\!&g^{\prime}n_{0}\\ g^{\prime}_{12}n_{0}&g^{\prime}_{12}n_{0}&g^{\prime}n_{0}&\!\!\epsilon^{0}_{k}+g^{\prime}n_{0}\!\end{array}\right),$ (17)
@@
Figure 1 shows excitation spectra $f_{\pm}$ versus dimensionless wave vector $k^{\prime}=\hbar k/\sqrt{2mg^{(0)}n_{0}}$ for various finite-range couplings $z_{\rho(\sigma)}$. We have checked that Eq. (19) is simplified into Ref. [56] for vanishing the finite-range interaction coupling constant $z_{\rho(\sigma)}=0$. The competition between contact ($y$) and finite-range ($z$) interactions distinctly modifies density and spin excitations in Figs. 1(a) and (b), respectively. In the phonon regime ($k\to 0$):
@@
Spinor $y\leq-1$, $z_{\rho(\sigma)}=0$ Quantum droplet [23]
$y\leq-1$, $z_{\rho(\sigma)}>0$ Finite-range quantum droplet [49, 50]
$0<y<1$, $z_{\rho(\sigma)}=0$ Spin-density separation [19]
$0<y<1$, $0\leq z_{\rho(\sigma)}\leq 1$ Spin-density separation (finite-range) This work
@@
(i) Single-component system ($y=0$): Finite-range interactions ($z_{\rho(\sigma)}\neq 0$) govern the Equation of state (EOS) [43, 44, 47, 45]. Positive $z_{\rho(\sigma)}>0$ induces a dimensional crossover [46, 48], while $z_{\rho(\sigma)}=0$ yields a distinct dimensional crossover [57, 58, 59].
@@
$z_{\rho(\sigma)}=0$: Quantum droplet phase for $y\leq-1$ (stabilized by beyond-mean-field effects [23]) and 1<y<0-1<y<0 [23]; spin-density separation for $0<y<1$ [19].
$z_{\rho(\sigma)}\neq 0$: Quantum droplets persist for $y\leq-1$ [50, 49]. For $0<y<1$ and $0\leq z_{\rho(\sigma)}\leq 1$, finite-range interactions modify spin-density separation (this work’s focus). The system transitions to a mixed bubble phase at $y>0$ [27] and phase separation at $y>1$ [60].
@@
Figure 2: (a) Scaling function $h(z_{\rho})$ in Eq. (25) for $z_{\rho}\equiv 4mn_{0}g_{\rho}^{(2)}/\hbar^{2}$ varied in $[0,0.7]$ and discrete values of $z_{\sigma}\equiv 4mn_{0}g_{\sigma}^{(2)}/\hbar^{2}$. (b) Scaling function $I_{\text{QD}}(z_{\sigma})$ in Eq. (29) for $z_{\sigma}\equiv 4mn_{0}g_{\sigma}^{(2)}/\hbar^{2}$ varied in $[0,0.7]$ and discrete values of $z_{\rho}$.
III GROUND-STATE ENERGY AND QUANTUM DEPLETION IN $d$-DIMENSIONAL SYSTEM